# AI-Driven Grading and Moderation for Collaborative Projects in Computer Science Education

### v1

@@ Abstract
Collaborative group projects are integral to computer science education, as they foster teamwork, problem-solving skills, and industry-relevant competencies. However, assessing individual contributions within group settings has long been a challenge. Traditional assessment strategies, such as the equal distribution of grades or subjective peer assessments, often fall short in terms of fairness, objectivity, and scalability, particularly in large classrooms. This paper introduces a semi-automated, AI-assisted grading system that evaluates both project quality and individual effort using repository mining, communication analytics, and machine learning models. The system comprises modules for project evaluation, contribution analysis, and grade computation, integrating seamlessly with platforms like GitHub. A pilot deployment in a senior-level course demonstrated high alignment with instructor assessments, increased student satisfaction, and reduced instructor grading effort. We conclude by discussing implementation considerations, ethical implications, and proposed enhancements to broaden applicability.
@@ I Introduction
Group projects are foundational to computer science curricula, offering students hands-on experience in software engineering practices such as Agile methodologies, version control, collaborative coding, and iterative testing. These collaborative endeavors mirror professional environments and are invaluable in preparing students for real-world development workflows. However, assigning fair individual grades remains difficult, particularly when group sizes are large or contributions are unevenly distributed.
@@ II Related Work
The broader field of Artificial Intelligence in Education (AIEd) applies machine learning, educational data mining, and analytics to support adaptive instruction, feedback, and monitoring of engagement [6]. However, automated grading systems face criticism due to a lack of transparency, limited adaptability, and reliance on historical patterns, which may disadvantage students who have atypical learning styles.
Finally, ethical considerations such as algorithmic fairness, data privacy, surveillance, and student consent are critical [8]. These must be carefully addressed before integrating AI into high-stakes assessments, such as course grades and other evaluations.
@@ III System Design and Architecture
In this paper, we propose an AI grading system comprising three core modules: the Project Quality Assessment Module (PQAM), the Individual Contribution Analyzer (ICA), and the Grading Engine (GE). We will illustrate each one and explain how they interact with each other in the following sections.
@@ III-A Project Quality Assessment Module (PQAM)
This module is a comprehensive evaluation component designed to assess both the technical and qualitative integrity of a group software project. It leverages a range of automated tools and AI models to ensure the project meets academic and industry standards across multiple dimensions. The module supports objective grading, encourages best practices, and reduces manual workload for instructors. This module contains five submodules: Code Quality, Testing Coverage, Documentation, Functionality, and Usability. The final output from this module is a general Project Quality Score (PQS), representing the overall quality of the group’s project. Each module, by default, holds the same weight. We can make weights configurable per project based on learning outcomes and project context.
@@ III-A1 Code Quality
Second, it utilizes Style Adherence to verify code compliance with specific programming language standards. For example, we can use Python tools such as flake8, pylint, and black to verify compliance with PEP8 standards. For JavaScript, tools like ESLint or Prettier ensure the code follows established best practices and organizational rules.
Third, it utilizes Code Duplication Detection to apply tools such as SonarQube, PMD, or jscpd, which detect copy-pasted blocks of code, indicating poor abstraction and reusability practices.
@@ III-A2 Testing Coverage
The Testing Coverage submodule is crucial to project reliability, as a robust testing practice evaluates how extensively tests cover the codebase. The available coverage tools for Python include pytest-cov and coverage.py, which can generate online reports, including branch and path coverage. For Java, JaCoCo or JUnit integration helps assess method and class-level coverage. For JavaScript, Istanbul/NYC or Jest produces detailed test coverage maps. Furthermore, beyond just counting lines covered, some AI models or advanced scripts can analyze whether edge cases, exception handling, and input validations are adequately tested to enhance the depth of coverage.
@@ III-A3 Documentation
The Documentation submodule reflects a team’s ability to communicate and maintain their software. This part uses NLP (Natural Language Processing) techniques to assess the quality and relevance of project documentation. For example, for readability and clarity, we can utilize NLP models such as BERT, spaCy, or GPT-based classifiers to evaluate the grammar, structure, and coherence of README files and manuals. For completeness checks, we can utilize heuristics and AI models to verify the inclusion of essential elements, such as installation steps, usage instructions, architecture overview, contributor guidelines, and licensing. For inline comments and docstrings, we can utilize parser tools like Docformatter, Doxygen, or Sphinx to assess whether functions/classes are documented, or perform semantic similarity analysis to check whether the comment aligns with the code logic or is superficial. For project Wikis or notebooks, such as Markdown or Jupyter Notebooks, Notebook-based documentation is evaluated for its technical depth, clarity of thought, and instructional value.
@@ III-A4 Functionality
The Functionality submodule is a direct evaluation of whether the software works as intended. We can enable an Automated Test Execution process, where predefined or student-supplied unit tests and integration tests are executed in sandboxed environments (e.g., Docker containers). The grading system compares actual outputs with expected outcomes and flags deviations. If students use tools like GitHub Actions, Travis CI, or GitLab CI, the system checks for passing build statuses and continuous testing practices. Moreover, feature validation can be conducted by using AI-based log analysis or semantic parsing of the project requirements to match the implemented functionality against the deliverables.
@@ III-A5 Usability
The Usability submodule assesses how intuitive, responsive, and accessible the system is, particularly if the project includes a GUI or web-based interface. Several major components can be used to evaluate code usability. For example, heuristic evaluation (optional/manual augmentation) can apply usability heuristics (e.g., Nielsen’s 10 principles) to detect issues like inconsistent navigation, lack of feedback, or poor error messaging. UI testing frameworks such as Selenium, Cypress, or Puppeteer can be used to run scripted tests across various devices and browsers. For accessibility checks, Lighthouse, Axe-Core, or Pa11y can be integrated to verify WCAG (Web Content Accessibility Guidelines) compliance. Color contrast, screen reader compatibility, and keyboard navigation are some of the key elements analyzed. For responsiveness analysis, UI components are tested across various screen resolutions to verify compatibility with both mobile and desktop devices.
@@ III-B Individual Contribution Analyzer (ICA)
This analyzer is a core module designed to objectively assess each student’s involvement in a collaborative group software project. Traditional group assessments often obscure individual effort, leading to potential unfairness in grading. ICA tackles this by analyzing multi-faceted signals from the development lifecycle to quantify and qualify each student’s contribution. ICA operates across four key dimensions: Commit History Analysis, Code Ownership, Issue Tracker Participation, and Code Review Evaluation.
@@ III-B1 Commit History Analysis
Commit History Analysis evaluates version control activity, focusing on meaningful code contributions rather than superficial edits. First, it filters trivial changes. AI heuristics and regular expression (regex) rules filter out non-substantive changes, such as whitespace adjustments, minor comment updates, or variable renaming. This prevents students from artificially inflating their contribution through “commit spam.” Second, for temporal and behavioral analysis, the system can observe patterns such as frequency of commits, time of day, and burst activity (indicating rushed contributions). Clustering algorithms (e.g., DBSCAN) can be used to distinguish between organic and last-minute participation. Third, quantitative metrics such as line additions, deletions, and net code changes are calculated. Furthermore, special weighting can be applied to the commits involving test cases, documentation, or refactors.
@@ III-B4 Code Review Evaluation
Code Review Evaluation is a key component of peer code review, which is critical for software quality and learning. ICA evaluates how constructively and frequently each student participates. The system can conduct the following analysis, if not all, to allow peer code review. The first one is Review Frequency, where the number of pull request reviews and comments per student is tracked, and the active participation in feedback loops (suggesting improvements, identifying flaws) is noted. The second one is NLP-based Depth Analysis, which utilizes NLP models (e.g., BERT, RoBERTa) to assess whether a review identifies logical issues, suggests improvements, or references standards or best practices. Here, the surface-level or generic comments receive lower weighting. The third one is Sentiment and Tone Analysis, which detects whether reviews are constructive, neutral, or overly critical/toxic. This encourages professionalism and positive collaboration.
@@ III-B5 Scoring & Visualization
Finally, the scoring and visualization of ICA adopts multimodal aggregation. ICA uses a weighted rubric to combine scores from the four components into a contribution index for each student. Customizable weights enable instructors to emphasize code, communication, or design. A visualization dashboard with contribution heatmaps, pie charts, and timelines helps visualize who contributed to which files, the evolution of each student’s work over time, and participation spikes (e.g., just before deadlines). Furthermore, anomaly detection can be integrated into ICA, where AI can flag potential freeloading or over-contributors, and team imbalance alerts can prompt early intervention by instructors.
@@ III-C Grading Engine (GE)
To ensure fair grading across teams of varying sizes and dynamics, ICA outputs could be normalized. For example, we can use Z-score Normalization or Min-Max Scaling, where each student’s raw contribution score is scaled within the context of their team to allow for comparison and mitigate the impact of team composition. Also, extremely low contributions (e.g., <10%) may trigger floor limits, and extremely high contributions may be capped to prevent “grade hogging” and support equity. Instructors can optionally add bonuses for exemplary contributors or penalties for under-participation based on team evaluations or reflections.
@@
A key goal of GE is to promote clarity and fairness in the grading process. It will provide breakdown reports for students that show the PQAM score, NCS, and final grade, along with possible visual aids for instructors, such as radar charts and contribution timelines. It also provides feedback prompts that can auto-generate personalized feedback summaries, for example, “Your code coverage and documentation were strong, but code review participation was low.”, or “Project quality was excellent, but individual contribution was significantly below team average.” All grading decisions and overrides are documented for transparency in the event of appeals or disputes.
The Grading Engine (GE) acts as the bridge between AI-driven analysis and practical grading outcomes. By combining group quality and individual contribution metrics with anomaly detection and transparency features, GE ensures that every student receives a fair, evidence-based grade—while still allowing instructors to exercise judgment when necessary.
@@ IV System Implementation and Analysis
To validate the effectiveness, fairness, and scalability of the AI-based grading system, a pilot deployment was conducted during a software engineering course at a university in fall 2024. The course consisted of 20 students divided into five teams, working on collaborative projects over 8 weeks. Projects adhered to Agile methodologies and required the use of GitHub for version control, issue tracking, and code submissions. This system was implemented in Python, utilizing frameworks such as Flask for RESTful API development and PostgreSQL for relational data persistence. The frontend dashboards were created using React.js, allowing instructors and students to interact with project metrics, grades, and feedback visually. The key libraries and tools include (1) Scikit-learn, TensorFlow–for building machine learning models (e.g., contribution prediction, NLP classification), (2) Radon, Pylint–static code analysis for metrics such as cyclomatic complexity and style compliance, (3) GitPython–used to interface with Git repositories for extracting commit history and file authorship data, and (4) SpaCy, NLTK–employed in the NLP pipeline to process README files, code comments, and issue tracker text for evaluating documentation and collaboration.
@@ IV-A2 Step 2: Data Extraction
Repositories were programmatically mined using the GitHub REST API and GraphQL endpoints, extracting commit history and diffs, issue tracker logs, pull requests, code reviews, README files, wikis, and inline comments.
@@ IV-A5 Step 5: Dashboard Review
Instructors accessed a secure web portal to view breakdowns of team performance and individual contribution, adjust or approve final grades, and export results to the LMS.
A system flowchart was generated (as shown in Fig. 3), illustrating modular interactions from submission through evaluation, grading, and review. The diagram helps stakeholders understand the end-to-end automation and where human judgment is involved.
@@ IV-C Pilot Results and Analysis
Three anonymized cases illustrate the system’s effectiveness in capturing real-world dynamics. First, contributions were evenly distributed (within ±5%), indicating that AI-assigned grades were uniform and required no manual overrides. Second, one team member authored 58% of the production code and conducted most of the issue resolution. The system elevated their individual score accordingly, with instructor validation. Third, one student inflated activity with frequent whitespace and comment-only commits. The ICA module, using code diff analysis, downweighted these and flagged the case, avoiding unfair grade inflation.
This pilot result highlights three key advantages of utilizing an AI grading system in collaborative coding projects. First, it demonstrates scalability. It seamlessly managed a class of 20 students and could scale to 200+ with minimal performance degradation. It also indicates objectivity as it replaced subjective perceptions with data-driven, consistent evaluation. Finally, it demonstrates transparency, as all students (based on an after-class survey) appreciated visibility into how grades were computed, thereby improving trust in the assessment. At the same time, it enhances the engagement incentive because awareness of tracked metrics motivates students to participate consistently.
However, we are facing particular challenges if we entirely rely on the system to grade. Essential activities, such as design ideation, presentation preparation, and team facilitation, were more strenuous to quantify objectively. Additionally, students who failed to use GitHub or issue trackers consistently had fewer contributions. Moreover, students with weaker English skills or non-native grammar patterns risked lower documentation scores. Ongoing calibration is needed to avoid bias.
@@ V Ethical Considerations
Fairness must be actively monitored to avoid unintended bias. NLP-based documentation analysis or behavioral metrics may disadvantage non-native speakers or students with atypical collaboration patterns. Regular audits using fairness metrics, anonymized data during manual reviews, and calibration across cohorts can help maintain equitable grading outcomes. Privacy and data governance are equally critical since repository histories, communication logs, and contribution records contain personally identifiable information. These data must be handled in compliance with institutional and legal standards such as FERPA and GDPR, supported by informed consent, restricted data access, and well-defined retention and deletion policies.
Finally, human oversight remains essential. The system is designed to augment, not replace, instructor judgment. Manual review tools and structured appeal processes ensure that students retain agency and that instructors can interpret algorithmic outcomes within their pedagogical context. Ethical integration depends on shared responsibility, where technology provides analytical consistency and educators ensure empathy and contextual understanding. By embedding transparency, fairness, privacy, and oversight into its design, the proposed system promotes responsible AI use in education. It reinforces the principles of trust and equity that underpin effective learning environments.

### v2

@@ Abstract
Collaborative group projects are integral to computer science education, as they foster teamwork, problem-solving skills, and industry-relevant competencies. However, assessing individual contributions in group settings has long been challenging. Traditional assessment strategies, such as the equal distribution of grades or subjective peer assessments, often fall short in terms of fairness, objectivity, and scalability, particularly in large classrooms. This paper introduces a semi-automated, AI-assisted grading system that evaluates both project quality and individual effort using repository mining, communication analytics, and machine learning models. The system comprises modules for project evaluation, contribution analysis, and grade computation, and integrates seamlessly with platforms such as GitHub. A pilot deployment in a senior-level course demonstrated high alignment with instructor assessments, increased student satisfaction, and reduced instructor grading effort. We conclude by discussing implementation considerations, ethical implications, and proposed enhancements to broaden applicability.
@@ I Introduction
Group projects are foundational to computer science curricula, providing students with hands-on experience in software engineering practices, including Agile methodologies, version control, collaborative coding, and iterative testing. These collaborative endeavors mirror professional environments and are invaluable in preparing students for real-world development workflows. However, assigning fair individual grades remains difficult, particularly when group sizes are large or contributions are unevenly distributed.
@@ II Related Work
The broader field of Artificial Intelligence in Education (AIEd) applies machine learning, educational data mining, and analytics to support adaptive instruction, feedback, and engagement monitoring [6]. However, automated grading systems are criticized for a lack of transparency, limited adaptability, and reliance on historical patterns, which may disadvantage students with atypical learning styles.
Finally, ethical considerations, including algorithmic fairness, data privacy, surveillance, and student consent, are critical [8]. These must be carefully addressed before integrating AI into high-stakes assessments, such as course grades and other evaluations.
@@ III System Design and Architecture
In this paper, we propose an AI grading system comprising three core modules: the Project Quality Assessment Module (PQAM), the Individual Contribution Analyzer (ICA), and the Grading Engine (GE). We will illustrate each one and explain how they interact in the following sections.
@@ III-A Project Quality Assessment Module (PQAM)
This module is a comprehensive evaluation component that assesses both the technical and qualitative integrity of a group software project. It leverages a range of automated tools and AI models to ensure the project meets academic and industry standards across multiple dimensions. The module supports objective grading, promotes best practices, and reduces instructors’ manual workload. This module contains five submodules: Code Quality, Testing Coverage, Documentation, Functionality, and Usability. The final output of this module is a general Project Quality Score (PQS) that represents the overall quality of the group’s project. Each module, by default, holds the same weight. We can make weights configurable per project based on learning outcomes and project context.
@@ III-A1 Code Quality
Second, it uses Style Adherence to verify code compliance with specific programming-language standards. For example, we can use Python tools such as flake8, pylint, and black to verify compliance with PEP8 standards. For JavaScript, tools such as ESLint and Prettier ensure that code adheres to established best practices and organizational standards.
Third, it uses Code Duplication Detection to apply tools such as SonarQube, PMD, or jscpd, which detect copy-pasted code blocks, indicating poor abstraction and reusability practices.
@@ III-A2 Testing Coverage
The Testing Coverage submodule is crucial for project reliability, as robust testing practices assess the extent to which tests cover the codebase. The available coverage tools for Python include pytest-cov and coverage.py, which can generate online reports, including branch and path coverage. For Java, JaCoCo or JUnit integration helps assess method and class-level coverage. For JavaScript, Istanbul/NYC, or Jest produces detailed test coverage maps. Furthermore, beyond simply counting lines covered, some AI models or advanced scripts can analyze whether edge cases, exception handling, and input validation are adequately tested to enhance coverage depth.
@@ III-A3 Documentation
The Documentation submodule reflects a team’s ability to communicate and maintain their software. This part uses NLP (Natural Language Processing) techniques to assess the quality and relevance of project documentation. For readability and clarity, we can use NLP models such as BERT, spaCy, or GPT-based classifiers to evaluate the grammar, structure, and coherence of README files and manuals. For completeness checks, we can utilize heuristics and AI models to verify the inclusion of essential elements, such as installation steps, usage instructions, architecture overview, contributor guidelines, and licensing. For inline comments and docstrings, we can use parser tools such as Docformatter, Doxygen, or Sphinx to assess whether functions/classes are documented, or perform semantic similarity analysis to determine whether the comment aligns with the code logic or is superficial. For project Wikis or notebooks, such as Markdown or Jupyter Notebooks, Notebook-based documentation is evaluated for its technical depth, clarity of thought, and instructional value.
@@ III-A4 Functionality
The Functionality submodule is a direct evaluation of whether the software works as intended. We can enable an Automated Test Execution process that executes predefined or student-supplied unit and integration tests in sandboxed environments (e.g., Docker containers). The grading system compares actual outputs with expected outcomes and flags deviations. If students use tools like GitHub Actions, Travis CI, or GitLab CI, the system checks for passing build statuses and continuous testing practices. Moreover, feature validation can be performed using AI-based log analysis or semantic parsing of the project requirements to verify that the implemented functionality aligns with the deliverables.
@@ III-A5 Usability
The Usability submodule assesses the system’s intuitiveness, responsiveness, and accessibility, particularly when the project includes a GUI or web-based interface. Several major components can be used to evaluate code usability. For example, heuristic evaluation (optional/manual augmentation) can apply usability heuristics (e.g., Nielsen’s 10 principles) to detect issues like inconsistent navigation, lack of feedback, or poor error messaging. UI testing frameworks such as Selenium, Cypress, or Puppeteer can be used to run scripted tests across various devices and browsers. For accessibility checks, Lighthouse, Axe-Core, or Pa11y can be integrated to verify WCAG (Web Content Accessibility Guidelines) compliance. Color contrast, screen reader compatibility, and keyboard navigation are some of the key elements analyzed. For responsiveness analysis, UI components are tested across a range of screen resolutions to verify compatibility with both mobile and desktop devices.
@@ III-B Individual Contribution Analyzer (ICA)
This analyzer is a core module that objectively assesses each student’s involvement in a collaborative group software project. Traditional group assessments often obscure individual effort, potentially leading to unfair grading. ICA tackles this by analyzing multi-faceted signals from the development lifecycle to quantify and qualify each student’s contribution. ICA operates across four key dimensions: Commit History Analysis, Code Ownership, Issue Tracker Participation, and Code Review Evaluation.
@@ III-B1 Commit History Analysis
Commit History Analysis evaluates version control activity, focusing on meaningful code contributions rather than superficial edits. First, it filters trivial changes. AI heuristics and regular expression (regex) rules filter out non-substantive changes, such as whitespace adjustments, minor comment updates, or variable renaming. This prevents students from artificially inflating their contribution through “commit spam.” Second, for temporal and behavioral analysis, the system can observe patterns such as commit frequency, time of day, and burst activity (indicating rushed contributions). Clustering algorithms (e.g., DBSCAN) can be used to distinguish between organic and last-minute participation. Third, quantitative metrics, including line additions, deletions, and net code changes, are computed. Furthermore, special weighting can be applied to the commits involving test cases, documentation, or refactors.
@@ III-B4 Code Review Evaluation
Code Review Evaluation is a key component of peer code review, which is critical for software quality and learning. ICA assesses the extent to which each student participates constructively and frequently. The system can conduct the following analysis, if not all, to allow peer code review. The first is Review Frequency, in which the number of pull request reviews and comments per student is tracked, and active participation in feedback loops (suggesting improvements, identifying flaws) is noted. The second one is NLP-based Depth Analysis, which utilizes NLP models (e.g., BERT, RoBERTa) to assess whether a review identifies logical issues, suggests improvements, or references standards or best practices. Here, the surface-level or generic comments receive lower weighting. The third one is Sentiment and Tone Analysis, which detects whether reviews are constructive, neutral, or overly critical/toxic. This encourages professionalism and positive collaboration.
@@ III-B5 Scoring & Visualization
Finally, the scoring and visualization of ICA adopts multimodal aggregation. ICA uses a weighted rubric to combine scores from the four components into a contribution index for each student. Customizable weights enable instructors to emphasize code, communication, or design. A visualization dashboard with contribution heatmaps, pie charts, and timelines helps visualize who contributed to which files, the evolution of each student’s work over time, and participation spikes (e.g., just before deadlines). Furthermore, anomaly detection can be integrated into ICA, enabling AI to flag potential freeloading or over-contributors and to issue team imbalance alerts, prompting early intervention by instructors.
@@ III-C Grading Engine (GE)
To ensure fair grading across teams of varying sizes and dynamics, ICA outputs could be normalized. For example, we can use Z-score Normalization or Min-Max Scaling, in which each student’s raw contribution score is scaled within their team to enable comparison and mitigate the impact of team composition. Also, extremely low contributions (e.g., <10%) may trigger floor limits, and extremely high contributions may be capped to prevent “grade hogging” and support equity. Instructors may optionally award bonuses to exemplary contributors or impose penalties for under-participation based on team evaluations or reflections.
@@
A key goal of GE is to promote clarity and fairness in the grading process. It will provide breakdown reports for students that show the PQAM score, NCS, and final grade, along with possible visual aids for instructors, such as radar charts and contribution timelines. It also provides feedback prompts that can auto-generate personalized feedback summaries, for example, “Your code coverage and documentation were strong, but code review participation was low,” or “Project quality was excellent, but individual contribution was significantly below team average.” All grading decisions and overrides are documented for transparency in the event of appeals or disputes.
The Grading Engine (GE) acts as the bridge between AI-driven analysis and practical grading outcomes. By combining group quality and individual contribution metrics with anomaly-detection and transparency features, GE ensures that every student receives a fair, evidence-based grade while still allowing instructors to exercise judgment when necessary.
@@ IV System Implementation and Analysis
To validate the effectiveness, fairness, and scalability of the AI-based grading system, a pilot deployment was conducted during a software engineering course at a university in fall 2024. The course consisted of 20 students, divided into five teams, who worked on collaborative projects over 8 weeks. Projects adhered to Agile methodologies and required the use of GitHub for version control, issue tracking, and code submissions. This system was implemented in Python, using Flask for RESTful API development and PostgreSQL for relational data persistence. The frontend dashboards were created using React.js, allowing instructors and students to interact with project metrics, grades, and feedback visually. The key libraries and tools include (1) Scikit-learn, TensorFlow–for building machine learning models (e.g., contribution prediction, NLP classification), (2) Radon, Pylint–static code analysis for metrics such as cyclomatic complexity and style compliance, (3) GitPython–used to interface with Git repositories for extracting commit history and file authorship data, and (4) SpaCy, NLTK–employed in the NLP pipeline to process README files, code comments, and issue tracker text for evaluating documentation and collaboration.
@@ IV-A2 Step 2: Data Extraction
Repositories were programmatically mined using the GitHub REST API and GraphQL endpoints to extract commit history and diffs, issue tracker logs, pull requests, code reviews, README files, wikis, and inline comments.
@@ IV-A5 Step 5: Dashboard Review
Instructors accessed a secure web portal to view performance breakdowns and individual contributions, adjust or approve final grades, and export results to the LMS.
A system flowchart (Fig. 3) was generated to illustrate the modular interactions from submission through evaluation, grading, and review. The diagram helps stakeholders understand the end-to-end automation and where human judgment is involved.
@@ IV-C Pilot Results and Analysis
Three anonymized cases illustrate the system’s effectiveness in capturing real-world dynamics. First, contributions were evenly distributed (within ±5%), indicating that AI-assigned grades were uniform and required no manual overrides. Second, one team member authored 58% of the production code and handled most issue resolution. The system elevated their individual score accordingly, with instructor validation. Third, one student inflated activity with frequent whitespace and comment-only commits. The ICA module, using code-diff analysis, downweighted these and flagged the case, thereby avoiding unfair grade inflation.
This pilot result highlights three key advantages of utilizing an AI grading system in collaborative coding projects. First, it demonstrates scalability. It seamlessly managed a class of 20 students and could scale to 200+ with minimal performance degradation. It also demonstrates objectivity by replacing subjective perceptions with data-driven, consistent evaluations. Finally, it proves transparency, as all students (based on an after-class survey) appreciated visibility into how grades were computed, thereby improving trust in the assessment. At the same time, it enhances the incentive to engage because awareness of tracked metrics motivates students to participate consistently.
However, we face particular challenges when we rely entirely on the system to grade. Essential activities, such as design ideation, presentation preparation, and team facilitation, were more strenuous to quantify objectively. Additionally, students who did not use GitHub or issue trackers consistently made fewer contributions. Moreover, students with weaker English skills or non-native grammar patterns were at higher risk of lower documentation scores. Ongoing calibration is needed to avoid bias.
@@ V Ethical Considerations
Fairness must be actively monitored to avoid unintended bias. NLP-based documentation analysis or behavioral metrics may disadvantage non-native speakers or students with atypical collaboration patterns. Regular audits using fairness metrics, anonymized data during manual reviews, and calibration across cohorts can help maintain equitable grading outcomes. Privacy and data governance are equally critical since repository histories, communication logs, and contribution records contain personally identifiable information. These data must be handled in compliance with institutional and legal standards, such as FERPA and GDPR. They must be supported by informed consent, restricted data access, and well-defined retention and deletion policies.
Finally, human oversight remains essential. The system is designed to augment, not replace, instructor judgment. Manual review tools and structured appeal processes ensure that students retain agency and that instructors can interpret algorithmic outcomes within their pedagogical context. Ethical integration depends on shared responsibility, in which technology provides analytical consistency, and educators ensure empathy and contextual understanding. By embedding transparency, fairness, privacy, and oversight into its design, the proposed system promotes responsible AI use in education. It reinforces the principles of trust and equity that underpin effective learning environments.